# 2Università degli Studi di Siena, Via Roma 56, Siena, Italy

### v1

@@ Abstract
Context. Young massive stellar clusters (YMSCs) have emerged as potential -ray sources, after the recent association of a dozen YMSCs with extended -ray emission. The large size of the detected halos, comparable to that of the wind-blown bubble expected around YMSCs, makes the -ray detection of individual YMSCs rather challenging. As a result, the emission from most of the Galactic YMSCs could be unresolved, thus contributing to the diffuse -ray radiation observed along the Galactic Plane.
Aims. In this study, we estimate the possible contribution to the Galactic diffuse -ray emission from a synthetic population of YMSCs, and we compare it with observations obtained with different experiments, from 1 GeV to hundreds of TeV, in two regions of the Galactic Plane.
@@
Results. We find that the -ray emission from unresolved YMSCs can significantly contribute to the observed Galactic diffuse flux, especially in the inner part of the Galaxy. The result is independent of the assumed particle transport, but an important role is played by Wolf-Rayet stars. The predicted -ray flux should be considered as a lower limit, given that our calculation neglects the contribution of supernovae exploding in YMSCs.
@@
The detection of large extended sources in -ray astronomy is often a challenging task due to the limited resolution of the telescopes, which prevents a clear disentanglement of multiple sources in crowded regions of the Galactic Plane. In addition, the background subtraction is often problematic due to the relatively small field of view of most current instruments. These observational limitations make the detection of individual YMSCs very difficult. At present, it is reasonable to expect that most of the emission coming from YMSCs is detected as a non-resolved contribution to the large-scale Galactic emission. Alongside, recent analysis of the Galactic diffuse -ray emission has highlighted the presence of an excess above a few GeV up to PeV energies which is possibly associated to emission from an unresolved population of sources (Zhang et al., 2023; Cao et al., 2023a).
@@
We here consider the IMF provided by Kroupa (2001). When sampling the IMF, we fix the minimum and maximum stellar masses that can be generated to M⊙ and M⊙. is related to the minimum theoretical mass to support nuclear burning (Carroll & Ostlie, 1996), while is the maximum observed stellar mass in clusters (Zinnecker & Yorke, 2007)111More recent works investigating the cluster R 136 infer a maximum stellar mass as large as (Crowther et al., 2010). However, we here adopt the smaller value of 150 M⊙ to provide a more conservative estimate for the gamma-ray emission..
@@
We do not consider any post-main-sequence evolution (with the exception of stars in the Wolf Rayet phase, that will be investigated separately below): namely, all the stars with are removed from the cluster. However, we neglect the energy released by the supernova explosions in the calculations of the CR power. In this sense, the production of CRs in YMSCs is underestimated, which implies that our estimate of the -ray flux from YMSCs is, in fact, a lower limit.
As an additional case of investigation, we build a mock stellar population including the presence of Wolf-Rayet (WR) stars. To this purpose, we assume that all stars with M⊙ undergo a WR phase for 0.3 Myr after (Rosslowe & Crowther, 2015), with the choice of the minimum WR mass motivated by the results of stellar evolution codes, showing that the minimum mass to develop a WR is in the range 22–37 M⊙ (Eldridge & Vink, 2006).
@@ 2.1 Modeling stellar winds
where and are the helium fraction and metallicity of WR stars respectively (both normalized to solar values). is the bolometric luminosity of the WR, which can be calculated using an ad-hoc empirical mass-luminosity relation (see Equation (3) by Schaerer & Maeder, 1992). Note that since depends on the WR mass, it can drastically change in time due to severe mass losses associated to the wind. We account for this dependence by considering the time evolution of the stellar mass in the evaluation of the mass loss rate, i.e. . Finally, the kinetic luminosity of the WR wind is calculated using Equation (3), but considering a constant wind speed of 2000 km s-1 (Crowther, 2007).
@@ 3 Generating synthetic population of Galactic YMSCs
which gives the number of clusters born at time per unit mass, time and surface. As commonly done in the literature, we here assume that the dependence of on space, time and mass, can be factorized, so that , where , and are, in order, the total cluster formation rate, the cluster initial mass function (CIMF) and the cluster spatial distribution. The total cluster formation rate and initial mass function can both be inferred from observations. In this regard, we rely on the work by Piskunov et al. (2018), based on the Milky Way Star Cluster Survey, where the cluster initial mass function is modeled as a broken power law: for clusters with M⊙, the cluster mass function is . We neglect SCs with mass smaller than because the integrated wind kinetic luminosity of those clusters is of the entire YMSC population. The maximum cluster mass is derived from observations (Piskunov et al., 2018), and is M⊙, corresponding to the mass of Westerlund 1. Notice that in generating the mock SC distribution, the number of stars is used in place of the cluster mass, see Appendix A. As we are interested in young and massive stellar clusters, we consider only SCs with age between and Myr, because the wind luminosity drops substantially at later times. After accounting for the effect of cluster evolution, the age distribution of stellar clusters is found to be approximately uniform in the last few tens of Myrs (Piskunov et al., 2018). This implies a constant cluster formation rate in the time interval we are interested in. This rate can be estimated from the star formation rate of local young stellar clusters within 1 kpc as measured by Bonatto & Bica (2011). Assuming the CIMF from Piskunov et al. (2018)(underfilling case) leads to an average local cluster formation rate of Myr-1 kpc-2 (Menchiari, 2023).
Concerning the spatial distribution of YMSCs in the Milky Way, we account for the spiral structure of the Galaxy and assume that star clusters follow the radial distribution of giant molecular clouds. Details are reported in Appendix B.
@@
The aforementioned ingredients allow us to build a synthetic SC population: we find a total number of 747 clusters with average values for the wind luminosity and the mass loss rate equal to erg s-1 and M⊙ yr-1, respectively.
@@ 4 Cosmic ray acceleration at the collective wind termination shock of YMSCs
When the average distance between stars in a cluster is less than the size of individual stellar wind TS, the winds from massive stars are expected to combine to create a collective cluster wind. This scenario is typical of young clusters as a direct consequence of the mass segregation at birth (see Menchiari, 2023, and references therein). As the wind material is shocked and heated, it expands adiabatically in the interstellar medium (ISM) generating large bubble-like structures similar to those observed close to single massive stars (Weaver et al., 1977). A model for particle acceleration in these systems was presented by Morlino et al. (2021) based on the idea that particles are accelerated at the collective wind TS via the diffusive shock acceleration mechanism, and subsequently escape from the acceleration site experiencing a combination of advection and diffusion within the hot bubble, until reaching its border. From there, CRs are free to escape in the unperturbed ISM.
@@
The parameters are determined by the energy dependence of the diffusion coefficient, which is ultimately related to the MHD turbulence spectrum. In this work, we will consider three different diffusion regimes: Kolmogorov, Kraichnan and Bohm-like diffusion. The coefficients relevant for each of these three diffusion regimes were derived in Menchiari et al. (2024) and are reported in Table 1. The parameter is the spectral injection slope of accelerated particles. We fix for all clusters . This value results from the modelling of the Cygnus OB2 emission performed by Menchiari et al. (2024), but a similar spectrum is also deduced from -ray measurements of other SCs (see e.g. Aharonian et al., 2019; Yang & Aharonian, 2017; Aharonian et al., 2022; Liu et al., 2022; Mitchell et al., 2024) for which an energy spectrum of 2.3-2.4 is found.
@@
where is the distance from Earth and is the average target density, which is expected to be close to the external ISM density, as a result of the expected fragmentation of the swept up shell (Blasi & Morlino, 2023). The differential cross section for -ray production, , is taken from Kafexhiu et al. (2014) and based on SYBILL.
Since, the -ray emissivity integrated along the line of sight is expected to depend on radius only weakly (see Menchiari et al., 2024), in order to speed up the calculation, we use, as a spatial model for the sources, a uniform disk of size equal to that of the projected wind bubble.
5 YMSC contribution to diffuse -ray emission
We now focus on two regions of the sky for which recent observations of the diffuse -ray emission are available in the literature, and compute the corresponding emission of our synthetic YMSCs. The first region (named ROI1), is defined for and , while the second region (named ROI2) is defined for , and . For both ROI1 and ROI2, flux measurements are provided by Fermi-LAT (Zhang et al., 2023) and LHAASO (Cao et al., 2023a). In ROI1 additional measurements of the diffuse flux are provided by the ARGO experiment222The ARGO diffuse flux is extracted from a region slightly smaller than ROI1 and defined as and . However, given the small differences in extension and position in the sky, we decided to include ARGO data in our work. (Bartoli et al., 2015)
@@ 5 YMSC contribution to diffuse -ray emission
To be consistent with this modus operandi, we also mask the simulated emission in the two ROIs following a similar method to that implemented in the LHAASO analysis. For the ROI1 we use three different masks, referred to as M, M, and M (see Appendix C for additional details):
@@
M removes the inner Galactic Plane (, ) and the Local Arm (disk centered on [, ]). This mask is designed to emulate the removal of Galactic sources, such as pulsar wind nebulae (PWNe) or supernova remnants (SNRs), which are not considered in our realization of the Milky Way and that are predominantly located along the Galactic Plane, within the first and fourth Galactic Quadrants.
@@
M masks all those positions of the sky where the significance of the emission from YMSCs at 100 TeV, evaluated as described in Appendix C, is larger than 5 times the diffuse emission. This mask is intended to remove all clusters whose emission should be bright enough to be detectable by LHAASO. We here assume that all the clusters detected by LHAASO should be also detected by Fermi-LAT. Hence, we do not mask nor remove the emission from YMSCs that should have been detected by Fermi-LAT but not by LHAASO. We will comment a posteriori the implications of this assumption.
In ROI2 we use only one mask, named M, that is defined using the same method as for M (see Appendix C).
@@
To account for the uncertainty associated with the stochastic nature of the generation of the synthetic population of YMSCs, we simulate 100 different realizations of the Milky Way and calculate, for each of them, the contribution to the diffuse -ray flux, adopting different masks and assumptions on the particle diffusion. Figures 1 and 2 show the YMSC contribution to the diffuse -ray emission in ROI1 and ROI2, respectively, for the three types of diffusion coefficient considered. Figure 1 shows the emission from YMSCs after application of each of the three masks. In all plots the dashed lines represent the median flux per energy bin over the different (100) realisations of the Milky Way. The shaded areas cover instead the range between the 25 and 75 percentile of the distribution, including (color filled) or neglecting (striped) the contribution from Wolf-Rayet stars.
@@
In both ROIs, the -ray flux due to unresolved YMSCs is negligible when considering only the wind power from main sequence stars (). On the contrary, when WR stars are added to the stellar populations () the contribution from YMSCs can account for a substantial fraction of the measured flux at energies TeV for the ROI1, especially for the Kolmogorov and the Kraichnan cases. To better visualize the relative importance of the emission from YMSCs, in the last column of Figure 1 and in Figure 2 we also show the total expected diffuse -ray flux () when the contribution from the bulk of diffuse CRs (i.e. the CR sea) is considered: , where is the flux from the CR sea (see Appendix D). Noticeably, for the Kraichnan case, the contribution from YMSCs is able to account for the extra component associated to non-resolved sources, while in the Kolmogorov case the emission from YMSCs can only explain the observed flux in the range TeV. On the other hand, in the Bohm case, the contribution from YMSCs to the diffuse emission is negligible. This is because, when Bohm diffusion is adopted, many more clusters have -ray emission above the LHAASO threshold and are hence removed by applying the mask M (see Appendix C).
@@
In ROI2 the contribution from YMSCs is on average negligible, regardless of the presence of WR stars. However, no solid conclusion can be drawn due to the large statistical uncertainties in this region. In fact, due to the low content of YMSCs in the outer Galaxy, some realisations have a much larger -ray flux. However, it is likely that the knowledge of the population of YMSCs in the outer region of the Galaxy is complete or quasi-complete. Consequently, one could consider to rely directly on observed clusters from Gaia (see, e.g. Celli et al., 2023; Mitchell et al., 2024), rather than using our approach.
It is worth mentioning that the diffuse emission below GeV is likely to be slightly overestimated, as we did not subtract the emission from those YMSCs that would have been detected by Fermi-LAT but not by LHAASO. However, we do not expect this choice to lead to a large error because the most powerful clusters are subtracted anyway. In any case, in this spectral region the contribution of YMSCs is negligible with respect to that of the CR sea due to their harder spectrum.

### v2

@@ Abstract
Context. Young massive stellar clusters (YMSCs) have emerged as potential -ray sources after the recent association of a dozen YMSCs with extended -ray emission. The large size of the detected halos, comparable to that of the wind-blown bubble expected around YMSCs, makes the -ray detection of individual YMSCs rather challenging. As a result, the emission from most of the Galactic YMSCs could be unresolved, thus contributing to the diffuse -ray radiation observed along the Galactic Plane.
Aims. In this study, we estimate the possible contribution to the Galactic diffuse -ray emission from a synthetic population of YMSCs, and we compare it with observations obtained with different experiments, from 1 GeV to hundreds of TeV, in three regions of the Galactic Plane.
@@
Results. We find that the -ray emission from unresolved YMSCs can significantly contribute to the observed Galactic diffuse flux, especially in the inner part of the Galaxy, and that an important role is played by kinetic power injected by the Wolf-Rayet stellar winds. The predicted -ray flux should be considered as a lower limit, given that our calculation does not include the contribution of supernovae exploding in YMSCs.
@@
The detection of large extended sources in -ray astronomy is often a challenging task due to the limited resolution of the telescopes, which prevents a clear disentanglement of multiple sources in crowded regions of the Galactic Plane. In addition, the background subtraction is often problematic due to the relatively small field of view of most current instruments. These observational limitations make the detection of individual YMSCs very difficult.
At present, it is reasonable to expect that most of the emission coming from YMSCs is detected as a non-resolved contribution to the large-scale Galactic emission. Alongside, recent analysis of the Galactic diffuse -ray emission has highlighted the presence of an excess above a few GeV up to PeV energies which is possibly associated to emission from an unresolved population of sources (Zhang et al., 2023; Cao et al., 2023a).
@@
While we are well aware that Supernova explosions are likely to provide a sizeable contribution to the cluster kinetic energy, comparable to, or even larger than, that of stellar winds, we do not to include CR acceleration by these sources in the current calculation. This choice deserves some comment. In the first place, aiming at estimating a lower limit to the contribution of YMSCs to the diffuse -ray emission, we decided not to include SNRs because these could only increase our estimate and increase its uncertainty, due to the lack of a well established description of CR production in Supernova Remnants (SNRs) expanding inside a star cluster. Our choice is justified a-posteriori by our finding that even without inclusion of SNRs, and hence in the most conservative case, YMSCs can provide a sizeable contribution to the diffuse -ray emission.
@@
We here consider the IMF provided by Kroupa (2001). When sampling the IMF, we fix the minimum and maximum stellar masses that can be generated to M⊙ and M⊙. is related to the minimum theoretical mass to support nuclear burning (Carroll & Ostlie, 1996), while is the maximum observed stellar mass in clusters (Zinnecker & Yorke, 2007)111More recent works investigating the cluster R 136 infer a maximum stellar mass as large as (Crowther et al., 2010). However, we here adopt the smaller value of 150 M⊙ to provide a more conservative estimate for the -ray emission..
@@
We do not consider any post-main-sequence evolution (with the exception of stars in the Wolf Rayet phase, that will be investigated separately below): namely, all the stars with are removed from the cluster. We remind that, as anticipated in Section 1, we neglect the energy released by the supernova explosions in the calculations of the CR power.
As an additional case of investigation, we build a mock stellar population including the presence of Wolf-Rayet (WR) stars. To this purpose, we assume that all stars with M⊙ undergo a WR phase for 0.3 Myr after (Rosslowe & Crowther, 2015), with the choice of the minimum WR mass motivated by the results of stellar evolution codes, showing that the minimum mass to develop a WR is in the range 22–37 M⊙ (Eldridge & Vink, 2006). Lower mass WR stars might in fact be born as a result of binary evolution. Those stars may provide a sizeable contribution to the wind power, but mostly at late times. In fact, this channel is important for stars in the 10-20 range which leave the main sequence at ages Myr, hence this effect is bound to have a minor impact for the clusters considered in this work. In any case, neglect of this additional contribution can only lead to underestimate the wind kinetic luminosity, and hence make our estimate of the YMSC contribution to the diffuse -ray background a more conservative lower limit.
@@ 2.1 Modeling stellar winds
where and are the helium fraction and metallicity of WR stars respectively (both normalized to solar values). is the bolometric luminosity of the WR, which can be calculated using an ad-hoc empirical mass-luminosity relation (see Equation (3) by Schaerer & Maeder, 1992). Note that since depends on the WR mass, it can drastically change in time due to severe mass losses associated to the wind. We account for this dependence by considering the time evolution of the stellar mass in the evaluation of the mass loss rate, i.e. . Finally, the kinetic luminosity of the WR wind is calculated using Equation (3), but considering a constant wind speed of 2000 km s-1. In general, the wind speed in WR stars is not constant and depends on the type of WR. For example, WC and WO stars have a terminal wind speed in the range km-1 s-1 (Niedzielski & Skorzynski, 2002; Crowther, 2007), while WN stars have wind velocities spanning km-1 s-1 (Crowther, 2007).
@@ 3 Generating synthetic population of Galactic YMSCs
which gives the number of clusters born at time per unit mass, time and surface. As commonly done in the literature, we here assume that the dependence of on space, time and mass, can be factorized, so that , where , and are, in order, the total cluster formation rate, the cluster initial mass function (CIMF) and the cluster spatial distribution.
3.1 Mass distribution of star clusters
@@
The CIMF can be inferred from the from observation of the local SC population. In this regard, we rely on the work by Piskunov et al. (2018), based on the Milky Way Star Cluster Survey, where the CIMF is modeled as a broken power law: for clusters with M⊙, the CIMF is . We neglect SCs with mass smaller than because the integrated wind kinetic luminosity of those clusters is of the entire YMSC population. The maximum cluster mass is derived from observations (Piskunov et al., 2018), and is M⊙, corresponding to the mass of Westerlund 1. Notice that in generating the mock SC distribution, the number of stars is used in place of the cluster mass, see Appendix A.
3.2 Age distribution of star clusters
Similarly to the CIMF, the SCs age distribution can be retrieved from the local cluster population. As we are interested in YMSCs, we consider only star clusters with age between and Myr, because the wind luminosity drops substantially at later times. This can be seen in Figure 1, where we show the wind power as a function of time for two YMSCs with masses of 103 and 104 M⊙. Since the wind power depends on the number of massive stars, we also show the range in which it can vary, calculated as the 10% and 90% percentile out of 100 possible samplings of the stellar IMF. After accounting for the effect of cluster evolution, the age distribution of stellar clusters is found to be approximately uniform in the last few tens of Myrs (Piskunov et al., 2018). This implies a constant cluster formation rate in the time interval we are interested in. This rate can be estimated from the star formation rate of local young stellar clusters within 1 kpc as measured by Bonatto & Bica (2011). Assuming the CIMF from Piskunov et al. (2018)(underfilling case) leads to an average local cluster formation rate of Myr-1 kpc-2 (Menchiari, 2023).
3.3 Spatial distribution of star clusters
1.
2.
| Parameters | Arm 1 | Arm 2 | Arm 3 | Arm 4 | Local Spur⋆ |
| Ri [kpc] | 3.27 | 4.29 | 3.58 | 3.98 | 8.16 |
| [∘] | 9.87 | 10.51 | 10.01 | 8.14 | 2.71 |
| [∘] | 38.5 | 189 | 215.2 | 320.1 | 50.6 |
On top of the spiral structure, we also take into account the structure of the innermost region. Here we consider the presence of the Galactic bar and the Near 3 kpc and Far 3 kpc arms. The first is modeled as an ellipse with semi-major axis of 3.3 kpc and an aspect ratio of 4:10 (Churchwell et al., 2009). The Near/Far-3-kpc Arms are modeled using an elliptical annulus, with a semi-major axis of 4.1 kpc, and an aspect ratio of 0.54 (a semi-minor axis of 2.2 kpc) (Green et al., 2011).
Spiral Arm to 4, if ;
Local spur if and ;
Spiral Arm or Near/Far-3-kpc Arms and Galactic Bar, if kpc;
Near/Far-3-kpc Arms and Galactic Bar, if kpc.
Once all clusters are associated to a structure, we proceed to perturb their positions according to a Gaussian distribution, following the fundamental concept that both Spiral Arms and the Near and Far structure possess an intrinsic thickness. The width of the spiral arm increases with galactocentric distance, so that the probability of finding a source displaced from the center of the arm by , is (Faucher-Giguère & Kaspi, 2006):
As for the clusters belonging to the Near and Far arms, we extract from a Gaussian probability distribution the Galactic x and y coordinates with a spread of kpc, such that the probability at returns a scattering compatible with the observed radial thickness of 0.5 kpc (Green et al., 2011). Notice that the thickness of a spiral arm in YMSCs may be slightly different from that inferred for pulsars, however the result on the diffuse emission depends very weakly on this parameter.
At last, we generate the vertical coordinate () following the observed gas distribution, i.e. an exponential profile with a characteristic spread of 100 pc (Strong et al., 2000):
(see Appendix B).
@@ 3.3 Spatial distribution of star clusters
With the cluster formation rate defined it is possible to finally build the synthetic SC population. We find a total number of 747 clusters with average values for the wind luminosity and the mass loss rate equal to erg s-1 and M⊙ yr-1, respectively. Figure 2 shows a possible realization of the Milky Way YMSCs population.
@@ 4 Cosmic ray acceleration at the collective wind termination shock of YMSCs
When the average distance between stars in a cluster is less than the size of individual stellar wind TS, the winds from massive stars are expected to combine to create a collective cluster wind. This scenario is favored in young clusters as a direct consequence of mass segregation at birth (Lada & Lada, 2003), namely the fact that the most massive stars (hence those producing the most powerful winds) forms at the very core of the SC. In our mock SC population, the star distribution is not simulated. We simply assume that all SCs are compact enough to generate a collective TS. As the wind material is shocked and heated, it expands adiabatically in the interstellar medium (ISM) generating large bubble-like structures similar to those observed close to single massive stars (Weaver et al., 1977). A model for particle acceleration in these systems was presented by Morlino et al. (2021) based on the idea that particles are accelerated at the collective wind TS via the diffusive shock acceleration mechanism, and subsequently escape from the acceleration site experiencing a combination of advection and diffusion within the hot bubble, until reaching its border. From there, CRs are free to escape in the unperturbed ISM.
@@
The parameters are determined by the energy dependence of the diffusion coefficient, which is ultimately related to the MHD turbulence spectrum. In this work, we will consider three different diffusion regimes: Kolmogorov, Kraichnan and Bohm-like diffusion. The coefficients relevant for each of these three diffusion regimes are reported in Table 2. This coefficients are derived in Menchiari et al. (2024) and are valid for any YMSCs. The parameter is the spectral injection slope of accelerated particles. We fix for all clusters . This value results from the modelling of the Cygnus OB2 emission performed by Menchiari et al. (2024), but a similar spectrum is also deduced from -ray measurements of other SCs (see e.g. Aharonian et al., 2019; Yang & Aharonian, 2017; Aharonian et al., 2022; Liu et al., 2022; Mitchell et al., 2024) for which an energy spectrum of 2.3-2.4 is found.
@@
where is the distance from Earth and is the average target density, which is expected to be close to the external ISM density, as a result of the expected fragmentation of the swept up shell (Blasi & Morlino, 2024). The differential cross section for -ray production, , is taken from Kafexhiu et al. (2014) and based on SYBILL.
Since, the -ray emissivity integrated along the line of sight is expected to depend on the projected angular radius only weakly (see Menchiari et al., 2024), in order to speed up the calculation, we use, as a spatial model for the sources, a uniform disk of size equal to that of the projected wind bubble.
Finally, we note that the -ray emission depends almost linearly on the bubble volume due to the swept up mass. Hence, a correct estimate of the bubble size is important. The adiabatic model by Weaver et al. (1977) used here overestimates the bubble size because it does not account for the ISM external pressure nor for the radiative losses. While the former has a very minor impact, the latter may be relevant at times Myr (Silich & Tenorio-Tagle, 2013). Such a timescale may be even reduced due to the external shell fragmentation which increases the contact surface between the hot shocked wind and the cold ISM (Lancaster et al., 2021). On the other hand, when also SN explosion are taken into account, the effect of cooling will be compensated by the injection of additional power. Including all those effects is a not trivial task because it requires a full time dependent calculation able to account for the full 3D hydro-dynamical evolution of the bubble. However, at the end of Section 5 we will evaluate the impact of a possible bubble size reduction by arbitrarily reducing its radius.
5 The contribution of YMSC to diffuse -ray emission
We now focus on two regions of the sky for which recent observations of the diffuse -ray emission are available in the literature, and compute the corresponding emission of our synthetic YMSCs. The first region (named ROI1), is defined for and , while the second region (named ROI2) is defined for , and . For both ROI1 and ROI2, flux measurements are provided by Fermi-LAT (Zhang et al., 2023) and LHAASO (Cao et al., 2023a). In ROI1 additional measurements of the diffuse flux are provided by the ARGO experiment222The ARGO diffuse flux is extracted from a region slightly smaller than ROI1 and defined as and . However, given the small differences in extension and position in the sky, we decided to include ARGO data in our work. (Bartoli et al., 2015).
5.1 Masking the contribution of resolved sources
@@
To be consistent with this modus operandi, we also mask the simulated emission in the two ROIs following a similar method to that implemented in the LHAASO analysis. To do so, we use three different masks, referred to as , , and (see Appendix B for additional details):
@@
removes the inner Galactic Plane (, ) and the Local Arm (disk centered on [, ]). This mask is designed to emulate the removal of Galactic sources, such as pulsar wind nebulae (PWNe) or supernova remnants (SNRs), which are not considered in our realization of the Milky Way and that are predominantly located along the Galactic Plane, within the first and fourth Galactic Quadrants.
@@
masks all those positions of the sky where the significance of the emission from YMSCs at 100 TeV, evaluated as described in Appendix B, is larger than 5 times the diffuse emission. This mask is intended to remove all clusters whose emission should be bright enough to be detectable by LHAASO. We here assume that all the clusters detected by LHAASO should be also detected by Fermi-LAT. Hence, we do not mask nor remove the emission from YMSCs that should have been detected by Fermi-LAT but not by LHAASO. We will comment a posteriori the implications of this assumption.
5.2 Comparison with the observed galactic diffuse -ray emission
@@
To account for the uncertainty associated with the stochastic nature of the generation of the synthetic population of YMSCs, we simulate 100 different realizations of the Milky Way and calculate, for each of them, the contribution to the diffuse -ray flux, adopting different masks and assumptions on the particle diffusion. Figure 3 shows the YMSCs contribution to the diffuse -ray emission in ROI1 (top row) and ROI2 (bottom row) for the three types of diffusion coefficient considered. The emission from YMSCs is calculated after applying in both ROI1 and ROI2. In all plots the dashed lines represent the median flux per energy bin over the different (100) realisations of the Milky Way. The shaded areas cover instead the range between the 25 and 75 percentile of the distribution, including (color filled) or neglecting (striped) the contribution from WR stars.
@@
In both ROIs, the -ray flux due to unresolved YMSCs is negligible when considering only the wind power from main sequence stars (). On the contrary, when WR stars are added to the stellar populations () the contribution from YMSCs can account in some cases for a substantial fraction of the measured flux. To better visualize the relative importance of the emission from YMSCs, we also show the total expected diffuse -ray flux () when the contribution from the bulk of diffuse CRs (i.e. the CR sea) is considered: , where is the flux from the CR sea (see Appendix C).
More precisely, in ROI1, the contribution of YMSCs in the Kolmogorov case can explain the diffuse -ray emission from 10 GeV to 1 TeV. In the Kraichnan scenario, however, the contribution of YMSCs is higher and the total -ray emission exceeds that observed from 10 GeV to 10 TeV by 20%. This implies that in the Kraichnan case, the contribution from YMSCs is able to totally account for the excess observed in the diffuse galactic emission if the product between and is decreased by a factor 0.8. On the other hand, in the Bohm case, the contribution from YMSCs to the diffuse emission is negligible. This is because, when Bohm diffusion is adopted, many more clusters have -ray emission above the LHAASO threshold and are hence removed by applying the mask (see Appendix B).
@@
In ROI2 the contribution from YMSCs is on average negligible, regardless of the presence of WR stars, with the exception of the Kraichnan case, where YMSCs could significantly contribute at 100 GeV. However, no solid conclusion can be drawn due to the large statistical uncertainties in this region. In fact, due to the low content of YMSCs in the outer Galaxy, some realisations have a much larger -ray flux. However, it is likely that the knowledge of the population of YMSCs in the outer region of the Galaxy is complete or quasi-complete. Consequently, one could consider to rely directly on observed clusters from Gaia (see, e.g. Celli et al., 2023; Mitchell et al., 2024), rather than using our approach.
These results must be considered with a few caveats in mind. First, the diffuse emission below 100 GeV is likely to be slightly overestimated, as we did not subtract the emission from those YMSCs that would have been detected by Fermi-LAT but not by LHAASO. In any case, we do not expect this choice to lead to a large error because the most powerful clusters are subtracted anyway. Second, as mentioned in Section 4, because of radiation losses, the size of the bubble could be smaller, reducing the -ray emission of each cluster. On average a radius reduction between 20% and 50% may be expected. If we assume that every bubble is systematically reduced by 20% we found that the overall flux is reduced by roughly the same factor. On the other hand, for a reduction of 50%, the flux is reduced by a factor , and the emission from YMSCs is drastically suppressed. However, the latter case is somehow unrealistic for two reasons: first, radiation losses are in general expected to appear after 10 Myr (Silich & Tenorio-Tagle, 2013), and second, one should also consider the re-heating of the gas induced by supernova explosion in the cluster, which can compensate for the energy losses. Finally, we note that adding the contribution of SNe the total predicted flux would likely overshoot the data for the current values of CR acceleration efficiency and gas density, especially in the Kraichnan scenario. We expect that consistency with the data could still be recovered by adjusting the product between and , while still keeping both parameters in a reasonable range. However, fitting the observed spectrum is beyond the scope of this work, as it would also require the inclusion of additional populations of unresolved galactic -ray sources.
To summarize, we assess that non-resolved YMSCs do provide a considerable contribution to the diffuse galactic -ray emission in the inner region of the Galactic plane if the particle diffusion coefficient follows a Kolmogorov-like or Kraichnan-like spectral energy dependence. Such a conclusion is strengthened by the fact that our estimate neglects the emission coming from particles accelerated at SNRs.